OpenStudy (anonymous):

Write a linear system for: Tickets for a school play cost $4 for adults and $2 for students. At the end of the play, the school sold a total of 105 tickets and collected $360.

OpenStudy (anonymous):

Say x tickets of adults Therefore 105-x tickets of children 4x+2(105-x)=360

hero (hero):

x = adults y = kids x + y = 105 4x + 2y = 360

OpenStudy (anonymous):

Thank you.

hero (hero):

@NotSObright , it says to write a linear system

hero (hero):

@isaidrawr , you're welcome